Use summation notation to express the sum.
5+20+80+......+20,971,520

Answer by richwmiller(17219)

Tn = t * r^(n - 1)
20,971,520 = 5 * 4^(n - 1)
20,971,520 = 5 * 4^(n - 1)
4194304=4^(n - 1)
n=12
S=t*(1 - r^n)/(1 - r)
S=5*(1 - 4^12)/(1 - 4)
S = 27962025